ABC and CVP Analysis: Multiple Products

Good Scent, Inc., produces two colognes: Rose and Violet. Of the two, Rose is more popular. Data concerning the two products follow:

Rose

Violet

Expected sales (in cases)

51,000

10,200

Selling price per case

$103

$81

Direct labor hours

32,250

5,500

Machine hours

10,350

3,150

Receiving orders

50

27

Packing orders

102

55

Material cost per case

$50

$45

Direct labor cost per case

$9

$6

The company uses a conventional costing system and assigns overhead costs to products using direct labor hours. Annual overhead costs follow. They are classified as fixed or variable with respect to direct labor hours.

Fixed

Variable

Direct labor benefits

$          —  

$177,425

Machine costs

187,000*

237,825

Receiving department

210,500  

Packing department

101,000  

   Total costs

$498,500  

$415,250

* All depreciation

Required:

1. Using the conventional approach, compute the number of cases of Rose and the number of cases of Violet that must be sold for the company to break even. In your computations, round variable unit cost to the nearest cent and round the number of break-even packages to the nearest whole number.

Break-even cases of Rose

Break-even cases of Violet

2. Using an activity-based approach, compute the number of cases of each product that must be sold for the company to break even. In your computations, round all computed amounts to the nearest cent and round the number of break-even packages to the nearest whole number.

Break-even cases of Rose

Break-even cases of Violet
Answer #1
Please give positive ratings so I can keep answering. Thanks!
Good Scent, Inc.
Answer 1 Note
Allocation of costs on the basis of direct labor hours
Particulars Rose Violet Total
Direct labor hours        32,250.00        5,500.00      37,750.00 A
Total Variable overhead 415,250.00 B
Direct labor hour rate for Variable overhead             11.00 C=A/B
Variable overhead allocated      354,750.00     60,500.00 415,250.00 D=C*A
Expected sales (in cases)        51,000.00      10,200.00 E
Variable overhead per case                  6.96                5.93 F=D/E
Particulars Rose Violet Total
Direct labor hours        32,250.00        5,500.00      37,750.00 See A
Total Fixed overhead 498,500.00 G
Direct labor hour rate for Fixed overhead             13.21 H=G/A
Fixed overhead allocated      425,870.86     72,629.14 498,500.00 I=H*A
Contribution margin income statement
Particulars Rose Violet Total
Selling price per case              103.00             81.00
Material cost per case                50.00             45.00
Direct labor cost per case                  9.00                6.00
Variable overhead per case                  6.96                5.93 See F
Contribution per unit                37.04             24.07 J
Total Fixed overhead      425,870.86      72,629.14 498,500.00 See I
Breakeven (cases)        11,496.32        3,017.59     14,513.90 K=I/J
Answer 2 L M N O=M+N P=L/O
Calculation of ABC rates for variable overhead Cost Activity usage ABC rates
Activity cost pool Activity Measure Cost Rose Violet Total
Direct labor benefits Direct labor hours 177,425.00      32,250.00        5,500.00      37,750.00                4.70
Machine costs Machine hours 237,825.00      10,350.00        3,150.00      13,500.00             17.62
Total Cost assigned 415,250.00
See P See M Q=P*M
See N
R=P*N
Allocation table Rose Violet
Activity cost pool Activity Measure ABC Rates Activity used Cost assigned Activity used Cost assigned
Direct labor benefits Direct labor hours                4.70      32,250.00 151,575.00        5,500.00      25,850.00
Machine costs Machine hours             17.62      10,350.00 182,332.50        3,150.00      55,492.50
Total Cost assigned 333,907.50     81,342.50 S
Expected sales (in cases)      51,000.00      10,200.00 See E
Variable overhead per case                6.55                7.97 T=S/E
U V W X=V+W Y=U/X
Calculation of ABC rates for Fixed overhead Cost Activity usage ABC rates
Activity cost pool Activity Measure Cost Rose Violet Total
Machine costs Machine hours 187,000.00      10,350.00        3,150.00      13,500.00             13.85
Receiving department Receiving orders 210,500.00             50.00             27.00             77.00        2,733.77
Packing department Packing orders 101,000.00           102.00             55.00           157.00           643.31
Total Cost assigned 498,500.00
See Y See V Z=Y*V See W AA=Y*W
Allocation table Rose
Violet
Activity cost pool Activity Measure ABC Rates Activity used Cost assigned Activity used Cost assigned
Machine costs Machine hours             13.85      10,350.00 143,366.67        3,150.00      43,633.33
Receiving department Receiving orders        2,733.77             50.00 136,688.31             27.00      73,811.69
Packing department Packing orders           643.31           102.00      65,617.83             55.00      35,382.17
Total Cost assigned 345,672.81 152,827.19 AC
Contribution margin income statement
Particulars Rose Violet Total
Selling price per case              103.00             81.00
Material cost per case                50.00             45.00
Direct labor cost per case                  9.00                6.00
Variable overhead per case                  6.55                7.97 See T
Contribution per unit                37.45             22.03 AB
Total Fixed overhead      345,672.81 152,827.19 498,500.00 See AC
Breakeven (cases)          9,229.56        6,938.73     16,168.29 AD=AC/AB